# Cause and Effect: Can Large Language Models Truly Understand Causality?

### v1

@@ 1 Introduction
Given the growing reliance on AI systems to make consequential, mission-critical decisions, we need to enhance the causal reasoning capabilities of LLMs [26, 29] revealed significant limitations in LLMs’ causal reasoning capacities. While they may mimic causal language, most need a genuine comprehension of causal mechanisms. This is concerning as it could propagate misinformation or lead to unreliable predictions. Bridging this causal reasoning gap is an active area of research.
@@ 2 Related Work
Various approaches have been explored in the literature to understand and enhance causal reasoning with LLMs. For example, in this paper [29], the authors assess the ability of LLMs to answer causal questions while discussing their strengths and weaknesses. They discuss the potential of integrating explicit and implicit causal modules to enhance LLMs’ capabilities in causal reasoning. However, they do not provide a concrete architecture or implementation.
@@
We aim to enhance the effectiveness of all four aspects of causal reasoning in addition to what has been done in [31]. Our method will enhance causal reasoning by incorporating explicit knowledge from knowledge graphs such as ConceptNet.
[28] critically examined the capabilities of LLMs in causal reasoning and inference while [29] argue that although LLMs can mimic causal language, they lack genuine causal understanding, coining the term “causal parrots”.
Through meta-structural causal models, [14] reveal limitations in LLMs’ causal reasoning. They find that LLMs trained on code (Code-LLMs) outperform text-only models in abductive and counterfactual reasoning, highlighting the value of programming structure for causal abilities.
@@
There has been growing research that examines the symbolic capabilities[27] of LLMs. [17] challenges criticisms about LLMs’ lack of extended structure and meaning grounding, arguing more research is needed to understand their linguistic representations. Ji et al. [8] utilize causal graphs to map complex pathways between prompts and code generation, improving transparency.
Applications and limitations of LLMs are also analyzed. [5] survey diverse applications but highlight ethical concerns, biases, and resource demands. Comparing inductive reasoning between humans and LLMs, [6] find similarities and differences in specific reasoning areas.
Other work focuses on evaluation and security. [10] developed a benchmark for assessing LLMs’ causal reasoning through symbolic questions. [30] reveal overfitting risks from human feedback and neurons disproportionately influencing outputs, demonstrating LLMs’ attack vulnerabilities. [31] systematically categorizes LLM evaluations into knowledge, reasoning, reliability, and safety competencies.
@@ 3 Approach
CARE-CA Hybrid Causal LLM Framework: Our approach combines the explicit, structured causal reasoning of ConceptNet knowledge graphs coupled with counterfactual sentences to improvise the causal understanding of LLMs. This novel architecture aims to surpass traditional decoder or encoder-only models by leveraging the rich semantic knowledge base of ConceptNet with advanced contextual inference capabilities and ’alternate scenarios’ of the contextual sentences to further aid the LLMs in understanding the causality of scenarios.
We are proposing a novel approach called CARE-CA, designed to enhance AI systems’ causal reasoning capabilities. This methodology is particularly adept at integrating rich contextual information and exploring counterfactual scenarios to arrive at a more robust understanding of causality in complex problems.
@@
1. Contextual Knowledge Integrator (CKI): CKI enriches the AI’s reasoning process with relevant external knowledge from databases like ConceptNet, providing a deep contextual backdrop against which causal relationships can be examined.
@@ 4.1 Data
To develop and evaluate our CARE-CA framework, we employed six distinct datasets. Each dataset serves a specific function within our research, ranging from training the model’s causal reasoning capabilities to evaluating its performance in various causal reasoning tasks. All experiments were performed with a dataset split of 75%-25% for train test sets, and 3 runs were conducted for each dataset model combination. Dataset for Causal Reasoning Identification (CRI):
@@
Context A detailed narrative context provides the backdrop for each scenario. These narratives describe situations where multiple events or factors coincide, potentially influencing outcomes. The contexts are designed to be realistic and thought-provoking, setting the stage for causal reasoning and counterfactual exploration.
Causal Questions For each scenario, a set of causal questions is provided to challenge the models’ abilities in causal reasoning. These questions are categorized into two main types:
@@
Choices and Answers Each question is accompanied by a set of choices, one designated as the correct answer. For cause-effect questions, the choices represent potential influencing factors, while for counterfactual questions, the choices depict possible outcomes under different circumstances. The correct answers are carefully labeled to facilitate evaluation.
@@ 4.2 Experimental Details
4.2.1. Causal Relationship Identification:
@@
Procedure: CARE-CA parses narrative texts to pinpoint and articulate the causal connections between events or actions.
Evaluation Metrics: We measure performance using Accuracy to accurately capture the model’s effectiveness in accurately identifying causal relations.
4.2.2. Counterfactual Reasoning:
@@
Procedure: The framework generates counterfactual statements based on given scenarios, demonstrating its understanding of how different conditions might alter outcomes.
Evaluation Metrics: Accuracy in counterfactual generation and the Quality of Counterfactual Statements, assessed by human evaluators for depth and realism.
4.2.3. Causal Discovery:
@@
Procedure: CARE-CA analyzes texts to identify underlying causal mechanisms that are not explicitly mentioned, leveraging both contextual cues and counterfactual reasoning.
Evaluation Metrics: Discovery Accuracy to evaluate how often CARE-CA correctly identifies new causal links, and the Novelty of Discoveries to gauge the uniqueness and insightfulness of the uncovered causal relationships.
Implementation Notes: Across all tasks, CARE-CA integrates explicit knowledge from ConceptNet and what-if scenarios from counterfactual to enhance structured causal reasoning abilities. This combination allows CARE-CA to approach causal reasoning tasks with a level of sophistication and nuance that mirrors human cognitive processes, setting a new standard for AI in causal reasoning.
@@ 4.3 Results
Causal Relationship Identification: On the Cladder dataset, the CARE-CA model led with a standout performance, achieving a 63% accuracy, indicating its strong capability to identify causal relationships. Decoder models like BERT and RoBERTa also showed commendable efforts, with BERT achieving a 53% accuracy and RoBERTa at 50.3%, showcasing their utility in parsing complex narratives for causal reasoning, albeit with varying degrees of success across metrics. The decoder model T5 highlighted its proficiency with a balanced performance, showcasing the effectiveness of its decoding capabilities in causal reasoning tasks.
On the Com2sense dataset, the decoder models encountered diverse challenges, with CARE-CA again leading at 67.1% accuracy, suggesting its consistent ability to navigate causal reasoning tasks. Decoder models like T5 and Gemini Pro demonstrated significant adaptability, achieving accuracies of 65.4% and 65.8%, respectively, underscoring their potential in handling nuanced reasoning tasks.
@@
Counterfactual Reasoning: The time-travel dataset, focused on counterfactual reasoning, highlighted models’ challenges in understanding hypothetical scenarios. Roberta achieved the highest performance among the models tested, with an accuracy of 68.78%, potentially due to its robust contextual understanding capabilities. The Gemini Pro and Llama models scored 38.4% and 24.2%, respectively, suggesting that despite their extensive training data, they might struggle with tasks requiring deep counterfactual inference, underscoring the importance of specialized training or prompting for such tasks.T5 and GPT 3.5 models performed well with 61.7% and 63.2% respectively. Our method got a slight jump in accuracy from the best-performing decoders; however, due to information overload, it could not compete with relatively more straightforward encoders such as ALBERT with 68% accuracy.
@@ 5 Analysis
The observed performances underscore the complexity of causal reasoning tasks and the varying abilities of models to address them. The CARE-CA framework’s superior performance across several tasks suggests that its hybrid approach, which leverages explicit causal knowledge and counterfactual reasoning, significantly enhances causal inference capabilities. Models like BERT and RoBERTa exhibit strong foundational abilities in causal reasoning, likely benefiting from their diverse pre-training. However, tasks requiring nuanced understanding or domain-specific knowledge, such as counterfactual reasoning and causal explanation, highlight the limitations of general-purpose models and the value of specialized training or frameworks like CARE-CA.
@@
We also performed human evaluations for the COPA dataset on 100 samples. The evaluator was presented with examples where the CARE-CA framework and other leading LLMs (such as BERT, RoBERTa, and GPT-3.5) responded. The evaluators were tasked with rating the responses based on several criteria:
@@
The human evaluation also highlighted areas for improvement. Feedback from evaluators pointed to occasional challenges in handling highly domain-specific scenarios, especially in the e-care dataset, suggesting an avenue for further refining CARE-CA’s domain adaptation capabilities.
@@ A.0.2 Causal Explanation
We also experimented with Causal Explanation task, which lies at the core of understanding why things happen. rather than simply observing patterns or connections, which may not necessarily reveal causality, it delves deeper to pinpoint the direct cause-and-effect links between variables. This involves unraveling how certain factors or actions directly produce a specific outcome. Its significance spans across a wide range of disciplines, such as philosophy, science, social sciences, medicine, and engineering, as it enables us to grasp the intricate workings of complex systems and foresee the effects of altering certain variables.

### v2

@@ 1 Introduction
Given the growing reliance on AI systems to make consequential, mission-critical decisions, we need to enhance the causal reasoning capabilities of LLMs. [24, 26] revealed significant limitations in LLMs’ causal reasoning capacities. While they may mimic causal language, most need a genuine comprehension of causal mechanisms. This is concerning as it could propagate misinformation or lead to unreliable predictions. Bridging this causal reasoning gap is an active area of research.
@@ 2 Related Work
Various approaches have been explored in the literature to understand and enhance causal reasoning with LLMs. For example, in this paper [26], the authors assess the ability of LLMs to answer causal questions while discussing their strengths and weaknesses. They discuss the potential of integrating explicit and implicit causal modules to enhance LLMs’ capabilities in causal reasoning. However, this lacked a methodical implementation approach to accomplishing the same.
@@
[25] critically examines the capabilities of LLMs in causal reasoning and inference while [26] argues that although LLMs can mimic causal language, they lack genuine causal understanding, coining the term “causal parrots”.
[12] reveals the limitations in LLMs’ causal reasoning by leveraging meta-structural causal models. They find that LLMs trained on code (Code-LLMs) outperform text-only models in abductive and counterfactual reasoning, highlighting the value of programming structure for causal abilities.
Given the widespread implications of LLM causal reasoning capabilities, we aim to enhance the effectiveness of all four aspects of causal reasoning in addition to the LLM evaluation work done in [27]. Our method will specifically focus on enhancing the causal reasoning by incorporating explicit knowledge from knowledge graphs such as ConceptNet.
Notably, the study of causal reasoning has been a prominent focus of research in the field of natural language processing (NLP), especially with the emergence of large language models (LLMs). Prior research has evaluated these models’ capacities for causal reasoning, pointing out both their advantages and disadvantages.
One remarkable work is the CRAB benchmark proposed by [18], which evaluates the ability of LLMs to infer causal relationships between real-world events. The authors found that while LLMs can perform well on certain causal reasoning tasks, they struggle with more complex scenarios that require a deeper understanding of causality. Similarly, [8] investigated whether LLMs can infer causation from correlation, a crucial skill for causal reasoning. Their findings suggest that while LLMs can learn some causal patterns, they often fail to distinguish between causal and non-causal relationships, highlighting the need for more targeted approaches.
Additionally, [9] explored the impact of the causal direction of data collection on the performance of LLMs in causal reasoning tasks. They found that models trained on data with a specific causal direction perform better on tasks that align with that direction, underscoring the importance of dataset design in causal reasoning research. These studies provide a solid foundation for understanding the current state of causal reasoning in LLMs.
Prior research has explored various approaches to enhance the causal reasoning capabilities of LLMs. For example, [25] assessed the ability of LLMs to answer causal questions, discussing their strengths and weaknesses. The authors suggested the potential of integrating explicit and implicit causal modules to improve LLM performance, which is a key principle underlying our CARE-CA framework.
While various past works have demonstrated the superior performance of GPT-3.5 and Gemini Pro in certain causal reasoning tasks, their work did not provide a concrete architecture to enhance these capabilities. In contrast, our CARE-CA framework goes a step further by proposing a novel hybrid approach that combines explicit causal knowledge from resources like ConceptNet introduced by [20] with the implicit reasoning capabilities of LLMs.
Interestingly, CARE-CA aims to provide a more comprehensive and effective solution for tackling a wider array of causal reasoning tasks by incorporating counterfactual reasoning and contextual prompting. Unlike previous methods that either relied on explicit or implicit causal reasoning, CARE-CA’s unique integration of these two complementary approaches sets it apart, allowing for a more robust and flexible causal understanding. This distinction enables CARE-CA to potentially outperform existing techniques in tasks such as causal relationship identification, counterfactual reasoning, and causal discovery, as demonstrated in our experimental evaluation.
@@ 3 Approach
CARE-CA Hybrid Causal LLM Framework: Our approach combines the explicit, structured causal reasoning of ConceptNet knowledge graphs coupled with counterfactual sentences to improvise the causal understanding of LLMs. This novel architecture aims to surpass traditional decoder or encoder-only models by leveraging the rich semantic knowledge base of ConceptNet with advanced contextual inference capabilities and ‘alternate scenarios’ of the contextual sentences to further aid the LLMs in understanding the causality of scenarios. The combination of these two provides relevant contextual information for the LLMs to understand the causal reasoning in question. We carry out a single variable test comparing the performance (X and Y) on CARE-CA v/s without and compare performance with accuracy, recall, precision and F1 scores.
@@
1. Contextual Knowledge Integrator (CKI): CKI enriches the AI’s reasoning process with relevant external knowledge graph - ConceptNet, providing a deep contextual backdrop against which causal relationships can be examined.
@@ 4.1 Data
To develop and evaluate our CARE-CA framework, we employed six distinct datasets. Each dataset serves a specific function within our research, ranging from training the model’s causal reasoning capabilities to evaluating its performance in various causal reasoning tasks. All experiments were performed with a dataset split of 75%-25% for train test sets, and 3 runs were conducted for each dataset model combination. We evaluated 5 LLMs - GPT-3.5, Mistral 7b, Gemini Pro, Llama 2, T5 using 5 datasets- COPA, Timetravel, CLadder, Com2sense and e-care on causal reasoning tasks, then compared the same LLMs against our proposed method: CARE-CA’. Dataset for Causal Reasoning Identification (CRI):
@@
Context: A detailed narrative context provides the backdrop for each scenario. These narratives describe situations where multiple events or factors coincide, potentially influencing outcomes. The contexts are designed to be realistic and thought-provoking, setting the stage for causal reasoning and counterfactual exploration.
Causal Questions: For each scenario, a set of causal questions is provided to challenge the models’ abilities in causal reasoning. These questions are categorized into two main types:
@@
Choices and Answers: Each question is accompanied by a set of choices, one designated as the correct answer. For cause-effect questions, the choices represent potential influencing factors, while for counterfactual questions, the choices depict possible outcomes under different circumstances. The correct answers are carefully labeled to facilitate evaluation.
@@ 4.2 Experimental Details
4.2.1 Causal Relationship Identification
@@
4.2.2 Counterfactual Reasoning
@@
4.2.3 Causal Discovery
@@
4.2.4 Evaluation Metrics
We used Accuracy, Precision and Recall as evaluation metrics for all the expeirments.
@@ 4.3 Results
Causal Relationship Identification: On the Cladder dataset, the CARE-CA model led with a standout performance, achieving a 63% accuracy, indicating its strong capability to identify causal relationships. The decoder model T5 highlighted its proficiency with a balanced performance, showcasing the effectiveness of its decoding capabilities in causal reasoning tasks.
On the Com2sense dataset, the decoder models encountered diverse challenges, with CARE-CA again leading at 67.1% accuracy, suggesting its consistent ability to navigate causal reasoning tasks.
@@
Counterfactual Reasoning: The time-travel dataset, focused on counterfactual reasoning, highlighted models’ challenges in understanding hypothetical scenarios. The Gemini Pro and Llama models scored 38.4% and 24.2%, respectively, suggesting that despite their extensive training data, they might struggle with tasks requiring deep counterfactual inference, underscoring the importance of specialized training or prompting for such tasks.T5 and GPT 3.5 models performed well with 61.7% and 63.2% respectively. Our method got a slight jump in accuracy from the best-performing decoders; however, due to information overload, it could not compete with relatively more straightforward encoders such as ALBERT with 68% accuracy.
@@ 5 Analysis
The observed performances underscore the complexity of causal reasoning tasks and the varying abilities of models to address them. The CARE-CA framework’s superior performance across several tasks suggests that its hybrid approach, which leverages explicit causal knowledge and counterfactual reasoning, significantly enhances causal inference capabilities. LLMs exhibit strong foundational abilities in causal reasoning, likely benefiting from their diverse pre-training. However, tasks requiring nuanced understanding or domain-specific knowledge, such as counterfactual reasoning and causal explanation, highlight the limitations of LLMs and the value of specialized training or frameworks like CARE-CA.
@@
We also performed human evaluations for the COPA dataset on 100 samples. The evaluator was presented with examples where the CARE-CA framework and other leading LLMs (such as T5, GPT-3.5) responded. The evaluators were tasked with rating the responses based on several criteria:
@@
Feedback from evaluators pointed to occasional challenges in handling highly domain-specific scenarios, especially in the e-care dataset, suggesting an avenue for further refining CARE-CA’s domain adaptation capabilities.
@@ A.0.2 Causal Explanation
We also experimented with Causal Explanation task, which lies at the core of understanding why things happen. rather than simply observing patterns or connections, which may not necessarily reveal causality, it delves deeper to pinpoint the direct cause-and-effect links between variables. . Its significance spans across a wide range of disciplines, such as philosophy, science, social sciences, medicine, and engineering, as it enables us to grasp the intricate workings of complex systems and foresee the effects of altering certain variables.
@@
A.0.3 CausalNet Dataset Generation
The CausalNet dataset was generated using GPT-4. The way we created the prompt was using a few shot approach giving a few examples to gpt-4 to understand the causal nature of the sentences and generate prompts that are diverse.
We used the following prompt: Develop a dataset composed of entries that challenge and enhance machine learning models’ understanding of causal relationships and counterfactual reasoning across various domains. Each entry in the dataset should follow this structure: "Context": A detailed description of a scenario that outlines a complex situation involving causal relationships. "Questions": A set of questions focusing on (1) identifying causal effects within the context and (2) exploring counterfactual scenarios, with multiple-choice answers to infer the model’s reasoning capabilities.